OpenStudy (anonymous):

(Coordinate Geometry) Write the equation of the circle in standard form with center and radius given. Center (8, 0), radius = √3

OpenStudy (zzr0ck3r):

(8-x)^2 + (0-y)^2 = sqrt(3)^2 (8-x)^2 = y^2 = 3

OpenStudy (anonymous):

HOW DO U DO IT? CAN U EXPLAIN IT BRIEFLY???

OpenStudy (zzr0ck3r):

are you asking me or him sradha?

OpenStudy (anonymous):

U

OpenStudy (anonymous):

\[\frac{(x-8)^2}{3}+\frac{y^2}{3}=1\]

OpenStudy (zzr0ck3r):

(x_1 - x)^2 + (y_1-y)^2 = r^2 where (x_1,y_1) is the point givin

OpenStudy (zzr0ck3r):

it's just pythagorean theorem on a little crack:)

OpenStudy (anonymous):

OK